A playground merry-go-round of radius R = 2.2C m h
Answer #1

We can solve this by the conservation of the angular momentum and taking the child as a particle at the adge of the merry-go

The initial angular momentum is

L_i=I_o\omega_o

L_i=(265kg\cdot m^2)(8.0{\rm rev/min})

L_i=2120kg\cdot m^2\cdot{\rm rev/min}

Then when the child is on the merry-go the moment of inertia is the moment of inertia of the merry-go plus the child, which is I_c=mR^2

I=I_c+I_o

I=mR^2+I_o

I=(23kg)(2.20m)^2+(265kg\cdot m^2)

I=376kg\cdot m^2

Then the angular momentum is

L_f=I\omega_f

And finally by the conservation of the angular momentum

L_i=I\omega_f

\omega_f=\frac{L_i}{I}

\omega_f=\frac{2120\,kg\cdot m^2\cdot{\rm rev/min}}{376\,kg\cdot m^2}

\omega_f=5.64\,{\rm rev/min}